.2, Two point charges 30 micrometer and -40 micrometer are held at the origin and at x 0.72 m, respectively particle q = 42 micrometer is released from rest at x 0.28 m. if the initial acceleration of the particle he magnitude of 1 x 10^5 m/s, what is its mass?
1
Expert's answer
2020-02-24T10:01:57-0500

As per the question,

"q_1=30\mu C"

"q_2=-40\mu C"

"q_1" and "q_2" are at the origin and at "x=0.72m"

"q=42\mu C"

q is at the x= 0.28m

acceleration ="10^5m\/sec"

Hence, the net force on the charge q,

"F_{net}=\dfrac{q_1q}{4\pi \epsilon_o r_1^2}+\dfrac{q_2q}{4\pi \epsilon_o r_2^2}"


"F_{net}=\dfrac{9\times 30\times 42\times 10^{-12}\times 10^{9}}{0.28^2}+\dfrac{9\times 40\times 42\times 10^{-12}\times 10^{9}}{0.44^2}"

"F_{net}=144.6+78"

"F_{net}=222.6N"

let the mass is m,

So, "F=ma"

"m=\dfrac{222.6}{10^5}=222.6\times 10^{-5}kg"

or,

"m=2.226gm"